Derive equations for finding the following, during the titration of a weak base using a strong acid;

1. Initial pH

2. pH at any given point before equivalence point

3. pH at equivalence point

4. pH at any given point after equivalence point
Answer #1

let the weak base be BOH

BOH    ------------>    B+ + OH-

Kb   = [B+][OH-]/[BOH]

[OH-] = Kb[BOH]/[B+]

Kw = [OH-][H+]

[H+ ] =   Kw / [OH-] = Kw/Kb[BOH]/[B+]

pH = 14 - log(Kb[BOH]/[B+])

2) At any point before equivalence point

its a basic buffer solution

Kb   = [B+]salt[OH-]/[BOH]base

[OH-] = Kb[BOH]base/[B+]salt

        taking negative log

pOH = pkb + log([B+]salt/[BOH]base)

pH = 14 - pOH = 14- {pkb + log([B+]salt/[BOH]base)}

3) At equivalence ,

its a case of Salt Hydrolsis so

pH = 7 -1/2pkb - logc